OpenStudy (anonymous):

At an automobile factory, 1849 parts are made in 4 hours. What is the average rate at which parts are made per hour?

OpenStudy (***[isuru]***):

hi, according to the data in the problem 1849 parts ---> 4 hours x parts ---> 1 hour 1849/x = 4/1 x = 1849/4 = 462.25 so... for 1 hour 462.25 parts are completed... but the average value is wt u already completed so the average value is 462... parts per hour

OpenStudy (***[isuru]***):

the question asks the average value of parts not partial pars so u cant round off 462.25 to 463...

OpenStudy (anonymous):

Right, 462 should be the answer

OpenStudy (anonymous):

thanks!

OpenStudy (***[isuru]***):

@annessa_erin did u get it ?

OpenStudy (***[isuru]***):

u r welcome!!

OpenStudy (anonymous):

Yes I did!